|4x + 2| >= 8
>= I think thats how id put greater than or equal to?
I just need help on how id go about doing the problem.
Where would i start and what would i do?
Ive been stuck for about 10 mins.
And keep going in circles with trying to do the problem.
Thanks,
Kerry.
: |4x + 2| >= 8
>= I think thats how id put greater than or equal to?
I just need help on how id go about doing the problem.
Where would i start and what would i do?
Ive been stuck for about 10 mins.
And keep going in circles with trying to do the problem.
Thanks,
Kerry.

Answer by Fombitz(1756)

abs(4x+2) >= 8
Think about absolute value problems as two separate problems.
The first is the positive solution.
The second is the negative solution.
The absolute value of x is x when x>0 and -x when x<0.
Hence the two solutions.
First let's do the positive solution.
(4x+2)>=8
4x>=6
x>=6/4
x>=3/2
Next, let's do the negative solution.
-(4x+2)>=8
-4x-2>=8
-4x>=10
x<=-10/4 Note the change in sign direction after division by negative number.
x<=-5/2
That is the solution set.
x<=-5/2 and x>=3/2